PLEASE NEED ANSWERS ASAP & SHOW ALL WORK THANKS
1) Multiply the following binomials. Simplify where possible.
A) (3 – x)(12 – x)
b) (9ab – 6b)(14b + 3a)
2) Ian factored 12x3y3 – 24x2y3 into 6xy(2x2y2 – 4xy2).
a. Verify that Ian’s factorization is correct, using multiplication.
b) Ian’s factorization is not considered ‘complete’. Explain why.
3) The area of a rectangle is 2x2 – 7x – 15. Determine binomials that represent the length and width of this rectangle.
Answer:

Answer by t0hierry(194)

(3 - x) (2 - x) = 6 -3x -2x + x^2 = 6 - 5x + x^2
(9ab – 6b)(14b + 3a) = 9*14 ab^2 + 9*3 a^2b - 6*14b^2 - 6*3 ab
= 3(42 ab^2 + 9a^2b - 28b^2 -6 ab)
You can simplify Ian's expression further
12x2y3 (x-2y)